Bad gums may be evidence of a bad heart. Researchers have discovered that 85% of people who have suffered a heart attack had periodontal disease, an inflammation of the gums. Only 29% of healthy people have this disease.

a) Suppose that in a certain community, heart attacks are quite rare, occurring with only 10% probability. If someone has periodontal disease, what is the probability that he or she will have a heart attack?

b) If 40% of the people in a community will have a heart attack, what is the probability that a person with periodontal disease in this community will have a heart attack?


1
Expert's answer
2021-02-25T00:42:06-0500

Solution:

a) We will assume that the statistics above part a) apply to this community. Then let us denote H - heart attack, D - periodontal disease. We are given:


"P(D)=0.29, P(D|H)=0.85, P(H) = 0.1" ,


We need to calculate the probability of Heart Attack, given that person has Periodontal Disease.

Using the law of conditional probability:


"P(A)*P(B|A)=P(A|B)*P(B)"


We get:


"P(H|D)*P(D)=P(H)*P(D|H)," P(H|D) is what we need to find.


"P(H|D) = \frac{P(H)*P(D|H)}{P(D)} = \frac{0.1*0.85}{0.29} \approx 0.293" or "29.3\%"


b) Assuming the statistics above the a) are correct for this case. We are given the following:


"P(D) = 0.29, P(D|H)=0.85, P(H)=0.4" and we basically need to calculate the same probability as in a) but with different given values:


"P(H|D) = \frac{P(H)*P(D|H)}{P(D)}=\frac{0.4*0.85}{0.29}=1.17"


It can mean either that all people who have Periodontal Disease will have a heart attack or that statistics presented above a) is incorrect.


Answer:

a) "0.293", or "29.3\%"

b) 1 or statistics are incorrect